Biil is flying his helicopter to meet Julia in Eaton.If he flies at 90mph he will arrive on time.If he flies 10mph slower,he will be their three minutes late.How far away is Bill from Julia.
Answer by richwmiller(17219)

90t=80(t+3/60)
t=2/5 hour =24 minutes
90*2/5=36 miles